Suppose that the time required to complete a 1040R tax form is normally distributed with a mean of 110 minutes and a standard deviation of 10 minutes. What proportion of 1040R tax forms will be completed in at most 104 minutes? Round your answer to at least four decimal places.
Answer by stanbon(57250)

Suppose that the time required to complete a 1040R tax form is normally distributed with a mean of 110 minutes and a standard deviation of 10 minutes. What proportion of 1040R tax forms will be completed in at most 104 minutes? Round your answer to at least four decimal places.
---------------
z(104) = (104-110)/10 = -6/10 = -0.6
--------
P(x <= 104) = P(z <= -0.6) = normalcdf(-100,-0.6) = 0.2743
=============
Cheers,
Stan H.
=============